# Traversable wormholes inside anisotropic magnetized neutron stars: physical properties and potential observational imprints

### v1

@@ III.1 Einstein field equations
$S_{\text{GR}\phi\chi\mathcal{B}}$ $=$ $\int d^{4}x\sqrt{-g}\left[\frac{\mathcal{R}}{2\kappa^{2}}-\frac{1}{2}A(\phi,\chi)\partial_{\mu}\phi\partial^{\mu}\phi-B(\phi,\chi)\partial_{\mu}\phi\partial^{\mu}\chi\right$. (7)
$\left.-\frac{1}{2}C(\phi,\chi)\partial_{\mu}\chi\partial^{\mu}\chi-V(\phi,\chi)+\mathcal{L}_{\text{matter}}+\mathcal{L}_{\mathcal{B}}\right].$
Here $\mathcal{R}$ denotes the Ricci scalar, $\kappa^{2}=8\pi Gc^{-4}$, where $G$ is the universal constant of gravitation, and $c$ denotes the speed of light; $A(\phi,\chi)$, $B(\phi,\chi)$, and $C(\phi,\chi)$ are arbitrary functions; $V(\phi,\chi)$ denotes the scalar fields potential; and $\mathcal{L}_{\text{matter}}$ is the Lagrangian of matter, and $\mathcal{L}_{\mathcal{B}}$ is the Lagrangian of magnetic field. Throughout the paper, we use $G=c=1$.
By varying the action in Eq. (7) with respect to the metric $g_{\mu\nu}$, we can obtain
$0$ $=$ $\frac{1}{2\kappa^{2}}\left(-\mathcal{R}_{\mu\nu}+\frac{1}{2}g_{\mu\nu}\mathcal{R}\right)$ (8)
$+\frac{1}{2}g_{\mu\nu}\left[-\frac{1}{2}A(\phi,\chi)\partial_{\rho}\phi\partial^{\rho}\phi-B(\phi,\chi)\partial_{\rho}\phi\partial^{\rho}\chi-\frac{1}{2}C(\phi,\chi)\partial_{\rho}\chi\partial^{\rho}\chi-V(\phi,\chi)\right]$
$+\frac{1}{2}\left[A(\phi,\chi)\partial_{\mu}\phi\partial_{\nu}\phi+B(\phi,\chi)(\partial_{\mu}\phi\partial_{\nu}\chi+\partial_{\nu}\phi\partial_{\mu}\chi)+C(\phi,\chi)\partial_{\mu}\chi\partial_{\nu}\chi\right]$
$+\frac{1}{2}\left(T_{\text{matter}\,\mu\nu}+T_{\mathcal{B}\,\mu\nu}\right).$
Here, $T_{\text{matter}\,\mu\nu}$ denotes the EMT of the matter sector, while $T_{\mathcal{B}\,\mu\nu}$ represents the EMT of the magnetic field. The Greek indices run from 0 to 3.
Furthermore, by varying the action with respect to the fields $\phi$ and $\chi$, one obtains
$\frac{1}{2}A_{\phi}\partial_{\mu}\phi\,\partial^{\mu}\phi+A\nabla^{\nu}\partial_{\mu}\phi+A_{\chi}\partial_{\mu}\phi\,\partial^{\mu}\chi+\left(B_{\chi}-\frac{1}{2}C_{\phi}\right)\partial_{\mu}\chi\,\partial^{\mu}\chi+B\nabla^{\mu}\partial_{\mu}\chi-V_{\phi}=0,$ (9)
$\left(-\frac{1}{2}A_{\chi}+B_{\phi}\right)\partial_{\mu}\phi\,\partial^{\mu}\phi+B\nabla^{\mu}\partial_{\mu}\phi+\frac{1}{2}C_{\chi}\partial_{\mu}\chi\,\partial^{\mu}\chi+C\nabla^{\mu}\partial_{\mu}\chi+C_{\phi}\partial_{\mu}\phi\,\partial^{\mu}\chi-V_{\chi}=0.$ (10)
@@
III.2 Elimination of the ghosts
First, it is important to note that the elimination of ghosts presented in this subsection follows the procedure outlined in Ref. [38]. Now consider a metric of a general spherically symmetric and time-dependent spacetime, given by
$ds^{2}=$ $-e^{2\nu(t,r)}dt^{2}e^{2\lambda(t,r)}dr^{2}+$
@@ III.2 Elimination of the ghosts
A detailed justification that this ansatz does not lead to any loss of generality can be found in Refs. [73, 74, 75].
According to Ref. [38, 73, 74, 75, 76], the functions $A(\phi,\chi)$ and/or $C(\phi,\chi)$ are often negative, which implies that $\phi$ and/or $\chi$ behave as ghosts. To eliminate the ghosts, appropriate constraints must be introduced by adding Lagrange multipliers $\lambda_{\phi}$ and $\lambda_{\chi}$. This results in the addition of an additional action $S_{\lambda}$, which reads [38, 74]
$S_{\lambda}=\int d^{4}x\sqrt{-g}\big[$ $\lambda_{\phi}\big(e^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\big)$
$+\lambda_{x}\big(e^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\big)\big].$ (13)
@@
The fluctuations of $\phi$ and $\xi$ can be written as [38, 74, 75]
@@
which means that both $\delta\phi$ and $\delta\chi$ vanish in the whole spacetime if we set the initial condition $\delta\phi=0$ and the boundary condition $\delta\chi\rightarrow 0$ when $r\rightarrow\infty$, and we are left with $\delta\phi=\delta\chi=0$ [38, 74, 75].
@@
It is worth noting that, as pointed out in Ref. [38], even within the model defined by the modified action $S_{GR\phi\chi\mathcal{B}}+S_{\lambda}$, the choice $\lambda_{\phi}=\lambda_{\chi}=0$ consistently satisfies the field equations. Consequently, any solution of Eq. (8) derived from the original action (7) also remains a solution in the modified model with action $S_{GR\phi\chi\mathcal{B}}+S_{\lambda}$.
@@
It is also worth noting that the constraints given by Eqs. (14) and (15) generalize the mimetic constraint introduced in Ref. [72], where nondynamical dark matter effectively emerges. In the present model—consistent with the discussion in Ref. [38]—what effectively arises is not nondynamical dark matter, but rather nondynamical exotic matter like a phantom.
III.3 Energy conditions and traversability
@@
First, we have to highlight that the total EMT in our model generally consist of contributions from ordinary matter, the magnetic field, and the scalar fields. In GR, the NEC, WEC, SEC, and DEC impose the following sequence of inequalities on the total EMT $T_{\text{tot}\,\mu\nu}$ [38, 77]:
$T_{\text{tot}\>\mu\nu}\>k^{\mu}k^{\nu}\geq 0,\qquad T_{\text{tot}\>\mu\nu}\>V^{\mu}V^{\nu}\geq 0,\qquad\left(T_{\text{tot}\>\mu\nu}-\frac{1}{2}g_{\mu\nu}T_{\text{tot}}\right)V^{\mu}V^{\nu}\geq 0,$
$T_{\text{tot}\>\mu\nu}\>V^{\mu}V^{\nu}\geq 0,\qquad\text{and}\qquad T_{\text{tot}\>\mu\nu}\>V^{\nu}\qquad\text{is not spacelike,}$
for any null vector $k^{\mu}$, $g_{\mu\nu}k^{\mu}k^{\nu}=0$, and for any timelike vector $V^{\mu}$, $g_{\mu\nu}V^{\mu}V^{\nu}<0$. For the system that we are considering in this work, generally, we have $T_{\text{tot}\>\mu}^{\nu}=\text{diag}(-\rho_{\text{tot}};p_{r,\text{tot}};p_{t,\text{tot}};p_{t,\text{tot}})$. Now the energy conditions writes [38, 76]
NEC: $\quad\rho_{\text{tot}}+p_{r,\text{tot}}\geq 0,\qquad\rho_{\text{tot}}+p_{t,\text{tot}}\geq 0,$
WEC: $\quad\rho_{\text{tot}}+p_{r,\text{tot}}\geq 0,\qquad\rho_{\text{tot}}+p_{t,\text{tot}}\geq 0,\qquad\rho_{\text{tot}}\geq 0,$
SEC: $\quad\rho_{\text{tot}}+p_{r,\text{tot}}\geq 0,\qquad\rho_{\text{tot}}+p_{t,\text{tot}}\geq 0,\qquad\rho_{\text{tot}}+p_{r,\text{tot}}+2p_{t,\text{tot}}\geq 0,$
DEC: $\quad\rho_{\text{tot}}\geq 0,\qquad\rho_{\text{tot}}\geq|p_{r,\text{tot}}|,\qquad\rho_{\text{tot}}\geq|p_{t,\text{tot}}|.$
@@
To make the wormhole traversable, we need the violation of the NEC in the vicinity of a throat [78]. It implies the violation of all energy conditions [38]. Before analyzing the traversability of the wormhole, we must first determine the total energy density $\rho_{\text{tot}}$, the total radial pressure $p_{r,\text{tot}}$, and the total tangential pressure $p_{t,\text{tot}}$. Once again we have to highlight that these quantities consist of contributions from ordinary matter, the magnetic field, and the scalar fields. To obtain them, one must first derive the explicit expressions for the functions $A$, $B$, $C$, and $V$.
@@
The metric function $e^{2\lambda(r)}$ is given by
@@
where $r_{0}$ denotes the radius of the wormhole throat. Furthermore, for $r\geq R_{s}$, where $R_{s}$ is the stellar radius, the exterior solution satisfies
@@
Note that even for the magnetized configuration of NSs, Eq. (20) can still be used as a good approximation, since the pressure and energy density associated with the magnetic field in the exterior region of magnetars are negligible (for $B_{s}=10^{15}\,\mathrm{G}$, both the magnetic energy density and pressure are of the order of $10^{-5}\,\mathrm{MeV\,fm^{-3}}$). Therefore, the Schwarzschild exterior metric remains a valid approximation. A similar treatment is also employed in Refs. [63, 71].
@@
$\mathcal{T}_{\mu\nu}=\begin{pmatrix}-\left(\rho+\dfrac{\mathcal{B}^{2}}{8\pi}\right)g_{tt}&0&0&0\\[6.0pt] 0&\left(p_{r}+\dfrac{\mathcal{B}^{2}}{24\pi}\right)g_{rr}&0&0\\[6.0pt] 0&0&\left(p_{t}+\dfrac{\mathcal{B}^{2}}{24\pi}\right)g_{\theta\theta}&0\\[6.0pt] 0&0&0&\left(p_{t}+\dfrac{\mathcal{B}^{2}}{24\pi}\right)g_{\varphi\varphi}\end{pmatrix}$. (21)
Now we can define
$\varrho\equiv\left(\rho+\dfrac{\mathcal{B}^{2}}{8\pi}\right),\>\>\>\>P_{r}\equiv\left(p_{r}+\dfrac{\mathcal{B}^{2}}{24\pi}\right),\>\>\>\>P_{t}\equiv\left(p_{t}+\dfrac{\mathcal{B}^{2}}{24\pi}\right),$ (22)
where we still have $\sigma=p_{r}-p_{t}=P_{r}-P_{t}$.
With the current EMT in our hands, and by using Eq. (8), the $(t,t)$, $(r,r)$, $(\theta,\theta)$ components can be written as
$e^{2\nu}\left(-\frac{A}{2}e^{-2\nu}-\frac{C}{2}e^{-2\lambda}-V\right)+e^{2\nu}\varrho=\frac{e^{-2\lambda+2\nu}}{\kappa^{2}}\left(\frac{2\lambda^{\prime}}{r}+\frac{e^{2\lambda}-1}{r^{2}}\right),$ (23)
$e^{2\lambda}\left(\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}-V\right)+e^{2\lambda}P_{r}=\frac{1}{\kappa^{2}}\left(\frac{2\nu^{\prime}}{r}-\frac{e^{2\lambda}-1}{r^{2}}\right),$ (24)
$r^{2}\bigg(\frac{A}{2}e^{-2\nu}-\frac{C}{2}e^{-2\lambda}-V\bigg)+r^{2}P_{t}=\>$ $\frac{1}{\kappa^{2}}\big\{e^{-2\lambda}[r(\nu^{\prime}-\lambda^{\prime})+r^{2}\nu^{\prime\prime}+r^{2}(\nu^{\prime}-\lambda^{\prime})\nu^{\prime}]\big\}.$ (25)
With a little algebra, we obtain
$A$ $=\frac{e^{2\nu}}{\kappa^{2}}\left\{e^{-2\lambda}\left[\frac{\nu^{\prime}+\lambda^{\prime}}{r}+\nu^{\prime\prime}+(\nu^{\prime}-\lambda^{\prime})\nu^{\prime}+\frac{e^{2\lambda}-1}{r^{2}}\right]\right\}-e^{2\nu}(\varrho+P_{r}-\sigma),$ (26)
$C$ $=\frac{e^{2\lambda}}{\kappa^{2}}\left\{-e^{-2\lambda}\left[-\frac{\nu^{\prime}+\lambda^{\prime}}{r}+\nu^{\prime\prime}+(\nu^{\prime}-\lambda^{\prime})\nu^{\prime}+\frac{e^{2\lambda}-1}{r^{2}}\right]\right\}-e^{2\lambda}\sigma,$ (27)
$V$ $=\frac{e^{-2\lambda}}{\kappa^{2}}\left(\frac{\lambda^{\prime}-\nu^{\prime}}{r}+\frac{e^{2\lambda}-1}{r^{2}}\right)-\frac{1}{2}(\varrho-P_{r}).$ (28)
Note that the function $B$ appearing in Eq. (7) can be determined by allowing for a time-dependent metric: in that case the $(t,r)$ component of the field equations does not vanish but yields $B=\frac{2\dot{\lambda}}{\kappa^{2}r}$. Since in the present time-independent metric $\dot{\lambda}=0$, we therefore obtain $B=0$.
From Eqs. (23)-(25), we have
$\rho_{\text{tot}}=\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}+V+\varrho=0,$ (29)
$p_{r,\text{tot}}$ $=$ $\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}-V+P_{r}$ (30)
$=$ $\frac{2r(r-r_{0})\nu^{\prime}-r_{0}}{\kappa^{2}r^{3}},$
$p_{t,\text{tot}}$ $=$ $\frac{A}{2}e^{-2\nu}-\frac{C}{2}e^{-2\lambda}-V+P_{t}$ (31)
$=$ $[2r^{2}(r-r_{0})(\nu^{\prime\prime}+\nu^{\prime 2})+r(2r-r_{0})\nu^{\prime}$
$+r_{0}]\left(2\kappa^{2}r^{3}\right)^{-1}.$
It is interesting to note that neither the anisotropy nor the magnetic field appears in Eqs. (29)–(31). This implies that, if the NEC is indeed violated, these two factors do not affect the traversability.
@@
An important remark from Ref. [38] should be emphasized: the violation of the energy conditions usually induces instabilities in the configuration. For instance, the sound speed may exceed the speed of light. In the wormhole spacetime constructed from the two-scalar model used in this work, however, such instabilities do not arise. This is because the two scalar fields are nondynamical; they neither propagate nor fluctuate. Consequently, no sound mode is generated from oscillations of the effective fluid produced by the scalar fields, even though this effective fluid violates the energy conditions.
IV Non-magnetized configuration of WH+NS systems
From Eq. (10), for the WH+NS model considered here one obtains

### v3

@@ 3.1 Einstein field equations
3.2 Elimination of the ghosts
First, it is important to note that the elimination of ghosts presented in this subsection follows the procedure outlined in Ref. Nojiri2024 . Now consider a metric of a general spherically symmetric and time-dependent spacetime, given by
$ds^{2}=$ $-e^{2\nu(t,r)}dt^{2}+e^{2\lambda(t,r)}dr^{2}$
@@ 3.2 Elimination of the ghosts
A detailed justification that this ansatz does not lead to any loss of generality can be found in Refs. Nojiri2021 ; NojiriNashed2024 ; Nojiri2024NPB .
According to Ref. Nojiri2024 ; Nojiri2021 ; NojiriNashed2024 ; Nojiri2024NPB ; Nojiri2023 , the functions $A(\phi,\chi)$ and/or $C(\phi,\chi)$ are often negative, which implies that $\phi$ and/or $\chi$ behave as ghosts. To eliminate the ghosts, appropriate constraints must be introduced by adding Lagrange multipliers $\Lambda_{\phi}$ and $\Lambda_{\chi}$. This results in the addition of an additional action $S_{\Lambda}$, which reads Nojiri2024 ; NojiriNashed2024
${\color[rgb]{0,0,0}S_{\Lambda}}=\int d^{4}x\sqrt{-g}\big[$ ${\color[rgb]{0,0,0}\Lambda_{\phi}}\big(e^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\big)$
$+{\color[rgb]{0,0,0}\Lambda_{\chi}}\big(e^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\big)\big].$ (17)
@@
The fluctuations of $\phi$ and $\xi$ can be written as Nojiri2024 ; NojiriNashed2024 ; Nojiri2024NPB
@@
which means that both $\delta\phi$ and $\delta\chi$ vanish in the whole spacetime if we set the initial condition $\delta\phi=0$ and the boundary condition $\delta\chi\rightarrow 0$ when $r\rightarrow\infty$, and we are left with $\delta\phi=\delta\chi=0$ Nojiri2024 ; NojiriNashed2024 ; Nojiri2024NPB .
@@
It is worth noting that, as pointed out in Ref. Nojiri2024 , even within the model defined by the modified action $S_{GR\phi\chi\mathcal{B}}+{\color[rgb]{0,0,0}S_{\Lambda}}$, the choice $\color[rgb]{0,0,0}\Lambda_{\phi}=\Lambda_{\chi}=0$} consistently satisfies the field equations. Consequently, any solution of Eq. (LABEL:field) derived from the original action (LABEL:action) also remains a solution in the modified model with action $S_{GR\phi\chi\mathcal{B}}+{\color[rgb]{0,0,0}S_{\Lambda}}$.
@@
It is also worth noting that the constraints given by Eqs. (18) and (19) generalize the mimetic constraint introduced in Ref. Chamseddine2013 , where nondynamical dark matter effectively emerges. The review of mimetic gravity can be referred to Ref. Vagnozzi . In the present model—consistent with the discussion in Ref. Nojiri2024 —what effectively arises is not nondynamical dark matter, but rather nondynamical exotic matter like a phantom.
It is important to clarify the role of the auxiliary scalar fields $\chi$ and $\phi$ introduced in this construction. These fields are not intended to represent independent propagating degrees of freedom, but rather serve as auxiliary variables that parametrize the geometry and generate an effective stress–energy tensor through the structure of the action. In particular, once the ansatz $\chi=r$ and $\phi=t$ is imposed, the functions $A(\phi,\chi)$, $C(\phi,\chi)$, and the potential $V(\phi,\chi)$ become fixed and give rise to an effective exotic matter sector capable of supporting the wormhole geometry.
From a physical perspective, the exotic matter emerges as an effective source induced by the constrained auxiliary fields. As the construction remains free of ghost instabilities, the scalar fields should therefore be understood as non-propagating auxiliary fields with a physical sourcing role, rather than as conventional matter fields or pure coordinate labels.
3.3 Energy conditions and traversability
@@
First, we have to highlight that the total EMT in our model generally consist of contributions from ordinary matter, the magnetic field, and the scalar fields. In GR, the NEC, WEC, SEC, and DEC impose the following sequence of inequalities on the total EMT $T_{\text{tot}\,\mu\nu}$ Nojiri2024 ; Carroll2004 :
@@
To obtain the traversable wormhole, we need the violation of the NEC in the vicinity of a throat Zaslavskii2007 . It implies the violation of all energy conditions Nojiri2024 . Before analyzing the traversability of the wormhole, we must first determine the total energy density $\rho_{\text{tot}}$, the total radial pressure $p_{r,\text{tot}}$, and the total tangential pressure $p_{t,\text{tot}}$. Once again we have to highlight that these quantities consist of contributions from ordinary matter, the magnetic field, and the scalar fields. To obtain them, one must first derive the explicit expressions for the functions $A$, $B$, $C$, and $V$.
@@
We employ an ansatz
@@
where $r_{0}$ denotes the radius of the wormhole throat. Furthermore, for $r\geq R_{s}$, where $R_{s}$ is the surface radius, the exterior solution satisfies
@@
The ansatzes shown by Eqs. (23) and (24) have also been used in the ad hoc formalism within Ref. Nojiri2024 .
Note that even for the magnetized configuration of NSs, Eq. (24) can still be used as a good approximation, since the pressure and energy density associated with the magnetic field in the exterior region of magnetars are negligible (for $B_{s}=10^{15}\,\mathrm{G}$, both the magnetic energy density and pressure are of the order of $10^{-5}\,\mathrm{MeV\,fm^{-3}}$). Therefore, the exterior metric remains a valid approximation. A similar treatment is also employed in Refs. Pattersons2024a ; Mallick2014 .
For the WH+NS system, near the throat $r\sim r_{0}$, the metric function $e^{2\lambda(r)}$ behaves as Nojiri2024
$e^{2\lambda(r)}\sim\frac{r_{0}}{r-r_{0}}e^{2\lambda_{0}},$ (25)
where $\lambda_{0}$ is a constant. The radial coordinate $r$ can be redefined as Nojiri2024
$l=2\sqrt{r_{0}(r-r_{0})}.$ (26)
Near the throat, the metric becomes
$ds^{2}$ $\sim$ $-e^{2\nu(l)}dt^{2}+e^{2\lambda_{0}}dl^{2}$ (27)
$+r_{0}^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2}).$
From the ansatz given in Eq. (23), one immediately obtains $\lambda_{0}=0$. According to Ref. Nojiri2024 , although the new radial coordinate $l$ is defined to be positive, the metric with the new radial coordinate $l$ admits an analytic continuation to the region $l<0$. We assume that the region $l>0$ corresponds to our spacetime, while the region $l<0$ represents another spacetime, which is smoothly connected to ours through the wormhole.
@@
Note that the function $B$ appearing in Eq. (LABEL:action) can be determined by allowing for a time-dependent metric: in that case the $(t,r)$ component of the field equations does not vanish but yields $B=\frac{2\dot{\lambda}}{\kappa^{2}r}$. Since in the present time-independent metric $\dot{\lambda}=0$, we therefore obtain $B=0$.
From Eqs. (LABEL:tt)-(LABEL:thetatheta), we have
${\color[rgb]{0,0,0}\rho_{\text{tot}}=\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}+V+\varrho,}$ (36)
${\color[rgb]{0,0,0}p_{r,\text{tot}}=\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}-V+P_{r},}$ (37)
${\color[rgb]{0,0,0}p_{t,\text{tot}}=\frac{A}{2}e^{-2\nu}-\frac{C}{2}e^{-2\lambda}-V+P_{t}.}$ (38)
After substituting the expressions of $A$, $C$, and $V$ from Eqs. (LABEL:A)-(LABEL:V) into Eqs. (36)-(38) we obtain
${\color[rgb]{0,0,0}\rho_{\text{tot}}=0},$ (39)
${\color[rgb]{0,0,0}p_{r,\text{tot}}=\frac{2r(r-r_{0})\nu^{\prime}-r_{0}}{\kappa^{2}r^{3}},}$ (40)
${\color[rgb]{0,0,0}p_{t,\text{tot}}=\frac{2r^{2}(r-r_{0})(\nu^{\prime\prime}+\nu^{\prime 2})+r(2r-r_{0})\nu^{\prime}+r_{0}}{2\kappa^{2}r^{3}}.}$ (41)
It is important to emphasize that the absence of explicit anisotropic or magnetic-field terms in the final expressions for the energy density and pressures does not imply that anisotropy or the magnetic field are physically absent from the system. Rather, within the present effective description, the total energy density, the total radial pressure, and the total tangential pressure admit the same functional form in terms of the metric functions, independently of whether anisotropy and/or magnetic fields are included in the matter sector.
As a consequence, the traversability conditions are governed by the total EMT, and remain unchanged both in the presence and in the absence of anisotropy and magnetic fields. In this sense, anisotropy and the magnetic field do not explicitly modify the traversability criteria, although they contribute implicitly to the effective geometry through the construction of the metric.
@@
An important remark from Ref. Nojiri2024 should be emphasized: the violation of the energy conditions usually induces instabilities in the configuration. For instance, the sound speed may exceed the speed of light. In the wormhole spacetime constructed from the two-scalar model used in this work, however, such instabilities do not arise. This is because the two scalar fields are nondynamical; they neither propagate nor fluctuate. Consequently, no sound mode is generated from oscillations of the effective fluid produced by the scalar fields, even though this effective fluid violates the energy conditions.
4 Non-magnetized configuration of WH+NS systems
Before discussing the WH+NS configurations, we briefly clarify how the corresponding pure NS solutions are obtained. The NS configurations considered in this work are constructed using the standard TOV equations for a spherically symmetric relativistic fluid, without a wormhole throat. These solutions serve solely as reference configurations, allowing for a direct comparison between ordinary NSs and the WH+NS systems discussed below. No modification of the standard TOV formalism is assumed for the pure NS case.
From Eq. (LABEL:fieldscalar), for the WH+NS model considered here, one obtains